OpenStudy (anonymous):

A car is traveling around a circular race track at 180 kmh-1 . If a single lap of the track is 2.4 km long, What is the angular velocity of the car in rads ? W= 180/ 2.4 = W= V/R

OpenStudy (anonymous):

Ignore the part that says W= 180/2.4 & w=V/R

OpenStudy (tkhunny):

\(\dfrac{180\;km}{hr}\cdot\dfrac{1\;rev}{2.4\;km}\cdot \dfrac{2\pi\;rad}{1\;rev} = ??\;\dfrac{rad}{hr}\) The infinite variety of the number one is astounding!